WAC Notes July 25 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes July 24 2019 << >> WAC Notes July 26 2019

July 25 2019

Production Run List

Start Slug 13, IHWP OUT Flip Right, FFB on always

      • Since we are using bcm_dg_ds, the width in the bcm is ~1.5X of what we used to have with bcm_an_ds3
  • Good - run3622
  • Good - run3623
  • Good - run3624
  • Good - run3625
    • Mini-run 0 had too large detector widths->Fixed after bad_event_cut
  • Good - run3626
  • Good - run3629
    • Aq mean was large for first two miniruns
    • bad_evt_cut applied, now good
  • Good - run3630ADDED from suspicious list
    • Detector widths are fluctuating between mini-runs.
    • First run with detector width started increasing
    • Applied bad_evt_cut, now good
  • Good - run3631ADDED from suspicious list
    • Detector widths become worse
    • Beam position change
    • Beam current fluctuating
    • Normalized detector yield is decreasing over time
  • Good - run3632ADDED from suspicious list
    • Regressed detector widths are even higher
    • Normalized detector yield is decreasing over time
    • Regressed detectors are correlated
  • Good - run3633ADDED from suspicious list
    • Normalized detector yield is increasing over time (opposite to run3632)
    • Regressed detectors are correlated
    • Regressed detector widths are larger
  • Good - run3634ADDED from suspicious list
    • Normalized detector yield is increasing over time (opposite to run3632)
    • Regressed detectors are correlated
    • Regressed detector widths are larger
  • Good - run3635ADDED from suspicious list
    • Normalized detector yield is increasing over time (opposite to run3632)
    • Regressed detectors are correlated
    • Regressed detector widths are larger
  • Good - run3636ADDED from suspicious list
    • Long beam trip during the run
    • Normalized detector yield is decreasing again
    • Regressed detector widths are larger
  • Good - run3640ADDED from suspicious list
    • This run was taken during collimator scan
    • Aq widths are too large
  • Good - run3641ADDED from suspicious list
    • Aq widths are too large (>200ppm)
    • Started with 1uA checking position then go to 70uA

Non-Production Runs and Bad Runs

Pedestal/Calibrations Updates

Issues

  • There were many pedestal and event cut problems. From now on, when an event cut isn't absolutely necessary it should be removed - see japan source code for reference of cuts here
  • The BCM appears to be falling apart. We are fixing it and running slug 12 on BCM DG DS in the meantime.
  • BPM 4ec had intermittant problems (iocse10 has issues) and so we are regressing with 4e again. BPM 11X is ok and stays in regression.


  • Increase in Detector widths and Collimator temperature HALOG COLLIMATOR HALOG
  • Increase in collimator temperature and raising neutron dose HALOG
  • The increase in detector widths started from late Swing Shift HALOG
  • Cavity BPM lock change HALOG

Other Status Updates